Rohan has some books and some tables. If he puts 9 books into each table he will have 2 tables left vacant. If he puts 6 books on each table he will have 3 books left.find the number of books?

Answer by Boreal(15235)

x is the number of tables
9(x-2)=6x+3; no remainder with first; remainder with second is the +3
9x-18=6x+3
3x=21
x=7 tables
9*5=45 books
6*7+3=45 books